OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the second and third equations, then the first and second equations. 5x + 2y + z = -2 3x + 4y + 3z = 2 -4x - 3y - 3z = 1

OpenStudy (anonymous):

3x + 4y + 3z = 2 -4x - 3y - 3z = 1 ---------------- -x +y +0 = 3 y = 3+x

OpenStudy (anonymous):

A. 5x + 2y = -2 3x + 4y = 2 B. -x + y = 3 -12x - 2y = 8 C. 5x + 2y = -2 -12x - 2y = 8 D. x - y = -3 3x + 4y = 2